A particle is projected from ground at some angle with the horizontal. let P be the point At maximum height H. At what height above the point P should the particle be aimed to have range equal to maximum height

Grade:11

1 Answers

Khimraj
5 years ago
max height = range
u2sin\Theta/2g  = 2u2sin\Thetacos\Theta/g
½ = 2cos\Theta
then
cos\Theta = ¼
then
tan\Theta = \sqrt{15}
let height above point is x then
tan\Theta  = (H+x)/(H/2) = \sqrt{15}
then x = (H/2)(\sqrt{15} – 2)
Hope it clears.